the equation is w = square root of (H/500). now I want to know the H value and how to expand this equation?
Thanks and regards
Samy

Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

The equation is w = square root of (H/500). now I want to know the H value and how to expand this equation?
-------
w = sqrt(H/500)
---
w^2 = H/500
---
H = 500*w^2
===============
Cheers,
Stan H.
===============

Answer by oberobic(2304)